# NeurIPS 2025 Workshop on Space in Vision, Language, and Embodied AI

### v1

@@ 4 Experiments
Datasets. In real-world scenarios, simultaneously capturing dynamic scenes with different camera effects is extremely challenging. Existing public datasets lack multi-view data for the same dynamic scene under different camera effects. To address this limitation, we constructed our own dataset using Blender 4.5 Blender Online Community (2025), which enabled us to: (1) simulate realistic physics, and (2) consistently reproduce identical dynamic scenes while applying various camera effects.
We constructed 8 distinct dynamic scenes across 4 indoor environments (basketball court, warehouse, living room and bathroom) using high-quality models from BlenderKit BlenderKit Community (2025) and other asset platforms. Each scene features diverse material properties and physical interactions, including bouncing balls, mechanical animations, and wind-influenced motions. For each scene, we established 50 camera viewpoints rendered under four different camera effects (pinhole, fisheye, rolling shutter, depth of field), generating 50-frame videos at 512×512 resolution. Camera placement and point cloud generation were automated using the PlenoBlenderNeRF plugin H. (2025); Raafat (2024), ensuring consistent geometric foundations across different camera effects. Detailed scene descriptions and technical implementation are provided in the supplementary material.
@@
HexPlane Cao and Johnson (2023) 23.1124 0.7956 0.2942 0.20
MSTH Wang et al. (2023a) 29.431 0.9023 0.1139 9.38
@@
HexPlane Cao and Johnson (2023) 18.3722 0.7343 0.5056 0.01
MSTH Wang et al. (2023a) 28.4692 0.9009 0.1540 0.57
@@
HexPlane Cao and Johnson (2023) 15.6489 0.6678 0.4142 21.6869 0.7480 0.2726 0.21
MSTH Wang et al. (2023a) 24.4222 0.8163 0.1764 26.7930 0.8571 0.1153 9.38
@@
HexPlane Cao and Johnson (2023) N/A 21.3468 0.7723 0.3162 0.21
MSTH Wang et al. (2023a) N/A 28.7003 0.8927 0.1190 9.35
@@
HexPlane Cao and Johnson (2023) 12 hours 238
MSTH Wang et al. (2023a) 8 mins 889
@@
Baselines. Our task of generating videos with camera effects from dynamic scenes is novel. For fair and meaningful evaluation, we compare against two state-of-the-art dynamic NeRF methods, HexPlane Cao and Johnson (2023) and MSTH Wang et al. (2023a), selected for their capabilities in 4D scene reconstruction and ray tracing. To ensure consistency, we apply the same camera effect rendering module used in our method to these baselines, eliminating differences in effect generation. We conduct both quantitative and qualitative comparisons.
@@
Evaluation Metrics. We employ a set of metrics to assess the visual quality of our rendered videos against the ground-truth data. Specifically, we use Peak Signal-to-Noise Ratio (PSNR) Huynh-Thu and Ghanbari (2008), Structural Similarity Index Measure (SSIM) Wang et al. (2004), and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018). We also report training time, inference speed (FPS), and storage for efficiency evaluation. For all metrics, we report the average score over all frames, camera views, and scenes. For the rolling-shutter rendering comparison, we evaluate our method with chunk sizes of $N_{c}=1$, $4$, and $16$.
@@
Ground Truth Hexplane Cao and Johnson (2023) MSTH Wang et al. (2023a) 4D-GRT(Ours) Ground Truth Hexplane Cao and Johnson (2023) MSTH Wang et al. (2023a) 4D-GRT(Ours)
@@
Qualitative comparison. For qualitative evaluation, we present the rendering results of 4D-GRT (ours), MSTH Wang et al. (2023a), and Hexplane Cao and Johnson (2023) in Fig. 4. Our method demonstrates higher-fidelity results compared to others. Specifically, the reconstruction quality of Hexplane Cao and Johnson (2023) is obviously worse than ours, and there are reconstruction artifacts and color temperature misalignment in the results of MSTH Wang et al. (2023a), which do not exist in our results.
Qualitative results on real-world dataset. We further validate on real-world dynamic scenes by reconstructing and rendering sequences from the Neural 3D Video (Plenoptic Video) dataset Li et al. (2022) with camera effects in Fig. 5; more qualitative examples are provided in the supplementary material. In addition, we present results as videos to better demonstrate temporal consistency, and also include rendering results on sequences from our own dataset in video form for visualization of scene dynamics.
@@
Figure 5: Qualitative results of 4D-GRT on the Neural 3D Video dataset  Li et al. (2022).
@@ Appendix A Dataset Construction
To fill this gap, we construct a synthetic multi-view dataset in Blender 4.5 Blender Online Community (2025). We create 8 dynamic indoor scenes across 4 environments (basketball court, warehouse, living room, bathroom), each with 50 viewpoints and 50 frames at $\times512$, rendered under four effects (normal, fisheye, rolling shutter, depth of field).
Section A.1 details the Blender Blender Online Community (2025) pipeline, tools, and plugins used for multi-view dynamic rendering, and Section A.2 describes the scenes and their material/physics characteristics.
@@ A.1 Scene Construction using Blender
To establish a dataset suitable for evaluating model performance under different camera effects, we selected Blender 4.5 Blender Online Community (2025) as our primary scene construction and rendering tool. Blender provides highly realistic physics simulation capabilities, enabling accurate modeling of object motion, collisions, bouncing, and other dynamic behaviors. Additionally, it ensures completely consistent scene configurations across multiple renders, guaranteeing perfect content alignment under different camera effects while avoiding the limitations of real-world filming where identical dynamics cannot be reproduced.
For scene sourcing, we selected 4 high-quality indoor scene models through the BlenderKit BlenderKit Community (2025) asset library. In addition, certain dynamic objects, such as the basketball model used in the basketball court scene, were also obtained from BlenderKit. In the scenes of living room and bathroom, the dynamic objects were obtained from other online asset platforms: the LEGO bulldozer model Heinzelnisse (2014) and the Box animation model RonKo (2023) from Blend Swap BlendSwap Community , the running cat model dailyfree3d (2019) and a pot plant model 3dhaupt (2019) from free3D.com Free3D . Within each indoor environment, we designed 2 different types of dynamic objects and animations, such as bouncing spheres, jumping cubes, or other wind-influenced motion behaviors. In total, we constructed 8 distinct dynamic scenes.
@@
Our 50 camera viewpoint placement, corresponding camera pose data, and initial scene point cloud generation were all accomplished through the PlenoBlenderNeRF H. (2025) Blender plugin that is originally forked from BlenderNeRF Raafat (2024). This plugin automatically distributes multiple cameras uniformly throughout the scene and generates corresponding initial point clouds from Blender mesh vertex data before rendering, ensuring complete consistency between rendering results under different camera effects and geometric foundations, facilitating subsequent quality analysis and comparison.
@@ Appendix B Video Generation Experiment
To evaluate the capabilities of state-of-the-art video generation models in synthesizing videos with realistic camera effects, we conduct comprehensive experiments using four leading commercial video generation models: OpenAI Sora OpenAI (2024), Google Veo 3 Google DeepMind (2025), Tencent HunyuanVideo Kong et al. (2025), and ByteDance Seedance 1.0 Mini Gao et al. (2025). Our evaluation focuses on three fundamental camera distortion effects: fisheye distortion, rolling shutter artifacts, and depth of field effects. These effects represent critical aspects of realistic video synthesis that require accurate understanding of camera optics and sensor behavior.
@@ C.1 Gaussian Deformation Network Architecture
We adopt the 4DGS Wu et al. (2024b) formulation of the Gaussian deformation network to predict the deformation of 3D Gaussians at a specific time step. Specifically, the network consists of a spatio-temporal structure encoder and a multi-head Gaussian deformation decoder.
@@ C.2 Training details
The hyperparameters in our experiments are mainly based on those reported in 4DGS Wu et al. (2024b). In our observation, we find the setting following 4DGS Wu et al. (2024b) will not perform well in scenes with rapidly moving objects. Therefore, we increase the learning rate, the capacity of the deformation field, the basic resolution of the plane modules and the training iterations. Specifically, we adopt the following setting for scenes with objects undergoing vigorous motion. The width of deformation MLP is set to $128$. The basic resolution of the plane modules is set to $128$, which is upsampled by $2$, 44 and $8$. The densification threshold is set as $2.5\times 10^{-5}$ and increases to $1\times 10^{-4}$. The training iteration in the coarse stage is set to $5\times 10^{4}$ and in the fine stage it is set to $8\times 10^{4}$.
We adopt a two-stage optimization strategy similar to 4DGS Wu et al. (2024b). In the first stage, we initialize the positions and colors of the 3D Gaussians using point clouds generated from Blender’s synthetic scenes. We then train the 3D Gaussian scene—without the deformation network—using only the first frame. This step reconstructs the static components of the scene and provides a strong initialization for dynamic objects. In the second stage, we introduce the deformation network and jointly optimize both the 3D Gaussians and the network across all time frames.
In 3DGS Kerbl et al. (2023), densification is guided by the magnitude of the 2D positional gradients of Gaussians. In contrast, since our rendering relies on ray tracing rather than rasterization, we follow the approach of Moenne-Loccoz et al. (2024a) and use the magnitude of the 3D gradient of the Gaussian means to decide when to densify a Gaussian. Additionally, as in Moenne-Loccoz et al. (2024a), we scale the gradient magnitude by the distance of the deformed Gaussian from the camera, encouraging densification in the background.

### v2

@@ 4 Experiments
Datasets. In real-world scenarios, simultaneously capturing dynamic scenes with different camera effects is extremely challenging. Existing public datasets lack multi-view data for the same dynamic scene under different camera effects. To address this limitation, we constructed our own dataset using Blender 4.5 [7], which enabled us to: (1) simulate realistic physics, and (2) consistently reproduce identical dynamic scenes while applying various camera effects.
We constructed 8 distinct dynamic scenes across 4 indoor environments (basketball court, warehouse, living room and bathroom) using high-quality models from BlenderKit [8] and other asset platforms. Each scene features diverse material properties and physical interactions, including bouncing balls, mechanical animations, and wind-influenced motions. For each scene, we established 50 camera viewpoints rendered under four different camera effects (pinhole, fisheye, rolling shutter, depth of field), generating 50-frame videos at 512×512 resolution. Camera placement and point cloud generation were automated using the PlenoBlenderNeRF plugin [29, 61], ensuring consistent geometric foundations across different camera effects. Detailed scene descriptions and technical implementation are provided in the supplementary material.
@@
HexPlane [10] 23.1124 0.7956 0.2942 0.20
MSTH [72] 29.431 0.9023 0.1139 9.38
@@
HexPlane [10] 18.3722 0.7343 0.5056 0.01
MSTH [72] 28.4692 0.9009 0.1540 0.57
@@
HexPlane [10] 15.6489 0.6678 0.4142 21.6869 0.7480 0.2726 0.21
MSTH [72] 24.4222 0.8163 0.1764 26.7930 0.8571 0.1153 9.38
@@
HexPlane [10] N/A 21.3468 0.7723 0.3162 0.21
MSTH [72] N/A 28.7003 0.8927 0.1190 9.35
@@
HexPlane [10] 12 hours 238
MSTH [72] 8 mins 889
@@
Baselines. Our task of generating videos with camera effects from dynamic scenes is novel. For fair and meaningful evaluation, we compare against two state-of-the-art dynamic NeRF methods, HexPlane [10] and MSTH [72], selected for their capabilities in 4D scene reconstruction and ray tracing. To ensure consistency, we apply the same camera effect rendering module used in our method to these baselines, eliminating differences in effect generation. We conduct both quantitative and qualitative comparisons.
@@
Evaluation Metrics. We employ a set of metrics to assess the visual quality of our rendered videos against the ground-truth data. Specifically, we use Peak Signal-to-Noise Ratio (PSNR) [38], Structural Similarity Index Measure (SSIM) [76], and Learned Perceptual Image Patch Similarity (LPIPS) [90]. We also report training time, inference speed (FPS), and storage for efficiency evaluation. For all metrics, we report the average score over all frames, camera views, and scenes. For the rolling-shutter rendering comparison, we evaluate our method with chunk sizes of $N_{c}=1$, $4$, and $16$.
@@
Ground Truth Hexplane [10] MSTH [72] 4D-GRT(Ours) Ground Truth Hexplane [10] MSTH [72] 4D-GRT(Ours)
@@
Qualitative comparison. For qualitative evaluation, we present the rendering results of 4D-GRT (ours), MSTH [72], and Hexplane [10] in Fig. 4. Our method demonstrates higher-fidelity results compared to others. Specifically, the reconstruction quality of Hexplane [10] is obviously worse than ours, and there are reconstruction artifacts and color temperature misalignment in the results of MSTH [72], which do not exist in our results.
Qualitative results on real-world dataset. We further validate on real-world dynamic scenes by reconstructing and rendering sequences from the Neural 3D Video (Plenoptic Video) dataset [44] with camera effects in Fig. 5; more qualitative examples are provided in the supplementary material. In addition, we present results as videos to better demonstrate temporal consistency, and also include rendering results on sequences from our own dataset in video form for visualization of scene dynamics.
@@
Figure 5: Qualitative results of 4D-GRT on the Neural 3D Video dataset  [44].
@@ 5 Conclusion and Limitations
Acknowledgement
This research is supported by National Science and Technology Council, Taiwan (R.O.C), under the grant number of NSTC-114-2221-E-001-016, NSTC-113-2634-F-002-008, NSTC-112-2222-E-A49-004-MY2, NSTC-113-2628-E-A49-023-, and Academia Sinica under the grant number of AS-CDA-110-M09 and AS-IAIA-114-M10.
@@ Appendix A Dataset Construction
To fill this gap, we construct a synthetic multi-view dataset in Blender 4.5 [7]. We create 8 dynamic indoor scenes across 4 environments (basketball court, warehouse, living room, bathroom), each with 50 viewpoints and 50 frames at $\times512$, rendered under four effects (normal, fisheye, rolling shutter, depth of field).
Section A.1 details the Blender [7] pipeline, tools, and plugins used for multi-view dynamic rendering, and Section A.2 describes the scenes and their material/physics characteristics.
@@ A.1 Scene Construction using Blender
To establish a dataset suitable for evaluating model performance under different camera effects, we selected Blender 4.5 [7] as our primary scene construction and rendering tool. Blender provides highly realistic physics simulation capabilities, enabling accurate modeling of object motion, collisions, bouncing, and other dynamic behaviors. Additionally, it ensures completely consistent scene configurations across multiple renders, guaranteeing perfect content alignment under different camera effects while avoiding the limitations of real-world filming where identical dynamics cannot be reproduced.
For scene sourcing, we selected 4 high-quality indoor scene models through the BlenderKit [8] asset library. In addition, certain dynamic objects, such as the basketball model used in the basketball court scene, were also obtained from BlenderKit. In the scenes of living room and bathroom, the dynamic objects were obtained from other online asset platforms: the LEGO bulldozer model [31] and the Box animation model [63] from Blend Swap [9], the running cat model [16] and a pot plant model [1] from free3D.com [22]. Within each indoor environment, we designed 2 different types of dynamic objects and animations, such as bouncing spheres, jumping cubes, or other wind-influenced motion behaviors. In total, we constructed 8 distinct dynamic scenes.
@@
Our 50 camera viewpoint placement, corresponding camera pose data, and initial scene point cloud generation were all accomplished through the PlenoBlenderNeRF [29] Blender plugin that is originally forked from BlenderNeRF [61]. This plugin automatically distributes multiple cameras uniformly throughout the scene and generates corresponding initial point clouds from Blender mesh vertex data before rendering, ensuring complete consistency between rendering results under different camera effects and geometric foundations, facilitating subsequent quality analysis and comparison.
@@ Appendix B Video Generation Experiment
To evaluate the capabilities of state-of-the-art video generation models in synthesizing videos with realistic camera effects, we conduct comprehensive experiments using four leading commercial video generation models: OpenAI Sora [57], Google Veo 3 [28], Tencent HunyuanVideo [41], and ByteDance Seedance 1.0 Mini [26]. Our evaluation focuses on three fundamental camera distortion effects: fisheye distortion, rolling shutter artifacts, and depth of field effects. These effects represent critical aspects of realistic video synthesis that require accurate understanding of camera optics and sensor behavior.
@@ C.1 Gaussian Deformation Network Architecture
We adopt the 4DGS [78] formulation of the Gaussian deformation network to predict the deformation of 3D Gaussians at a specific time step. Specifically, the network consists of a spatio-temporal structure encoder and a multi-head Gaussian deformation decoder.
@@ C.2 Training details
The hyperparameters in our experiments are mainly based on those reported in 4DGS [78]. In our observation, we find the setting following 4DGS [78] will not perform well in scenes with rapidly moving objects. Therefore, we increase the learning rate, the capacity of the deformation field, the basic resolution of the plane modules and the training iterations. Specifically, we adopt the following setting for scenes with objects undergoing vigorous motion. The width of deformation MLP is set to $128$. The basic resolution of the plane modules is set to $128$, which is upsampled by $2$, 44 and $8$. The densification threshold is set as $2.5\times 10^{-5}$ and increases to $1\times 10^{-4}$. The training iteration in the coarse stage is set to $5\times 10^{4}$ and in the fine stage it is set to $8\times 10^{4}$.
We adopt a two-stage optimization strategy similar to 4DGS [78]. In the first stage, we initialize the positions and colors of the 3D Gaussians using point clouds generated from Blender’s synthetic scenes. We then train the 3D Gaussian scene—without the deformation network—using only the first frame. This step reconstructs the static components of the scene and provides a strong initialization for dynamic objects. In the second stage, we introduce the deformation network and jointly optimize both the 3D Gaussians and the network across all time frames.
In 3DGS [40], densification is guided by the magnitude of the 2D positional gradients of Gaussians. In contrast, since our rendering relies on ray tracing rather than rasterization, we follow the approach of [52] and use the magnitude of the 3D gradient of the Gaussian means to decide when to densify a Gaussian. Additionally, as in [52], we scale the gradient magnitude by the distance of the deformed Gaussian from the camera, encouraging densification in the background.